OpenStudy (anonymous):

Find the x- and y- intercepts for this equation.. 2x - 4y =1

OpenStudy (lalaly):

for x intercept take y=0 2x=1 x=1/2 for y intercept take x=0 -4y=1 y=-1/4

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Do you know how you would graph this eqution ?

OpenStudy (lalaly):

u still ther?